OpenStudy (anonymous):

Medal for help :) The point (12,9) is included in a direct variation.What is the constant of varation ? a.1/2 b.3/4 c.1 d.2

OpenStudy (phi):

y = k x y/x = k can you find k using (12,9), which is an (x,y) pair?

OpenStudy (anonymous):

Ok going to try now

OpenStudy (acxbox22):

y=kx 9=k(12) 9/12=k/12 ??=k

OpenStudy (anonymous):

C

OpenStudy (phi):

how did you get 1 ?

OpenStudy (anonymous):

Ish now i'm getting 2

OpenStudy (phi):

show what you are doing

OpenStudy (acxbox22):